OpenStudy (anonymous):

Factor the ff. Trinomials 5x2 + 24x + 16 2x2 - 3x - 9 6x2 - 27x - 15 pls. answer, or give me some hints pls

OpenStudy (ash2326):

I'll help you with the first one \[5x^2 + 24x + 16\] A standard quadratic is of the form \[ax^2+bx+c\] here a=5, b=24 and c=16 first step: find \(a\times b\), it comes as \(16\times 5=80\) second step: find factors of a*c such that the sum of the two is b, here we need to find factors of 80 such that their sum is 24 , the two factors are 4 and 20 let's verify that 4*20=80 4+20=24 Do you get this?

OpenStudy (ash2326):

@michellepau ??

OpenStudy (anonymous):

oh wait ? where are you from ?

OpenStudy (ash2326):

yeah, take your time. Let me know when you need help

OpenStudy (anonymous):

my assignments is 1-0, can you answer, 1-10 pls ?

OpenStudy (ash2326):

@michellepau I'll explain you this one, then you try the others. That way you'll learn better:) Ok?

OpenStudy (anonymous):

okay .. thanks you :))

OpenStudy (ash2326):

You're welcome, did you understand till there?

OpenStudy (anonymous):

actually no, i understand if the language is pilipino or tagalog

OpenStudy (ash2326):

@michellepau Sorry I don't know pilipino or tagalog. Try to understand, it's just maths